Circle Circumfe Circle Diameter r is radius Finish A sprinter runs the curve of this 200 m in 10.34 s. Assume he ran in a lane which makes a semicircle (r 36.3 m) for the first part of the race. At 3.44 s into the race his speed is 4.0 m/s. At 8.0 s into the race his speed is 9.6 m/s. His speed after the curve was 12.8 m/s. What was the sprinters average tangential acceleration between 3.44 s and 8.0 s? Answer: Check
Answer #1

We know that the average tangential acceleration is

a_avg=(v2-v1)/(t2-t1)

That is

a_avg=(9.6-4.0)/(8-3.44) m/s2

a_avg=5.6/4.56 m/s2

a_avg=1.228. m/s2